# EC-LDA : Label Distribution Inference Attack against Federated Graph Learning with Embedding Compression

### v1

@@ 5 Our Methodology: EC-LDA
Discussion. Implementing EC-LDA requires GNN layers with a fully connected final layer and the use of the cross-entropy loss function, which makes EC-LDA particularly versatile within classification tasks on GNNs, such as node classification. Since the link prediction is a binary classification task, EC-LDA is still applicable to link prediction tasks as long as the last fully connected layer of the model has two output units and the cross-entropy loss function is used. For node classification, EC-LDA can attack the label proportions, while for link prediction, it can target the graph density.
6 Experiments
@@
  • RQ1 - How effective is EC-LDA with real-world graph datasets, specifically for tasks such as node classification and link prediction?
  • RQ2 - How robust is EC-LDA under different experimental variables?
  • RQ3 - How the effectiveness of EC-LDA changes in the defense of differential privacy?
6.1 Experimental Settings
Here, we introduce evaluation metrics, model architecture, baselines, and relevant to the experiments as follow. The details of datasets and hyper-parameter settings please refer to the Appendix E.
Evaluation Metrics: Inspired by Geng (2016), we use the following two evaluation metrics to fully demonstrate the effectiveness of EC-LDA: cosine similarity(cos-sim) and Jensen-Shannon divergence(JS-div). Cos-sim measures the similarity of two distributions and is applicable when the similarity of vectors is not directly related to the length of the vectors, while JS-div measures the distance between two distributions. Cos-sim takes values in the range $[-1,1]$. The larger the cos-sim, the closer the inferred label distribution is to the ground-truth label distribution, i.e., the more effective the attack is. The value of JS-div ranges from $[0,1]$, and the more effective the attack is, the smaller the JS-div is.
Baselines: We compare EC-LDA with three different attacks which are Infiltrator Meng et al. (2023), iLRG Ma et al. (2023), and LLG* Wainakh et al. (2021). Infiltrator infers the label of the victim node by adding a neighbor to the victim node and observing the output of the neighbor. It is worth noting that Infiltrator focuses on node-level attacks, and for comparison, we attack all training nodes with Infiltrator. LLG* and iLRG both reveal the number of each label, and we use the number of samples per label extracted by iLRG and LLG* to compute the label distribution of each client. Similar to EC-LDA, iLRG and LLG* also use the gradients of the last fully connected layer of the model, and when $E$ is greater than 1, we use $\Delta W^{(FC)}/E$ as an approximation of the gradients.
Dummy Training Data: Based on the previous analysis, we know that the smaller the variance of $I$, the better the attack performance. Since the dummy training data is also related to the variance of $I$ we generate it from a Gaussian distribution with a mean of 0 and a standard deviation of 0.001, containing 1000 nodes. This configuration enhances the attack performance.
@@ 6.1 Experimental Settings
Table 2: Performance of different attacks when training GCN, GAT, and GraphSAGE models on various datasets.
6.2 Attack Performance (RQ1)
@@
Node Classification. We evaluate the performance of EC-LDA with all datasets and all GNN types, which, for the node classification task, have a varying number of labels and are widely distributed, with 4 kinds of labels for nodes in the Facebook dataset and 70 kinds of labels for nodes in the CoraFull dataset, the datasets with the smallest and the largest number of classes, respectively. Table 2 shows the experimental results. From the comparison presented in Table 2, it is evident that EC-LDA consistently demonstrates exceptional performance across various datasets and three distinct GNN models, as indicated by the cos-sim scores consistently at or above 0.999 and the JS-div consistently at or below 0.002, aligning closely with the optimal values. Remarkably, EC-LDA’s outstanding performance remains consistent regardless of the number of labels, the number of nodes, and the specific GNN types, showcasing its broad applicability. Additionally, EC-LDA consistently outperforms other methods across all experiments.
Link Prediction. In the link prediction experiments, we set the same number of positive and negative edges. Please refer to Appendix B for the experimental results. In all experiments of link prediction, cos-sim and JS-div reached 1.000 and 0.000, respectively, demonstrating the stunning performance of EC-LDA.
6.3 Ablation Experiments (RQ2)
To explore the effectiveness of EC-LDA under varying parameters, we performed extensive ablation experiments on the node classification tasks. Specifically, we explored the impact of the number of GNN layers, $E$, $C$, and number of clients separately. Please refer to the Appendix D for relevant experiments regarding the number of clients.
Impact of the Number of GNN Layers. We show the effect of the number of GNN layers on the performance of EC-LDA in Figure 4, where we use three different network structures with 1, 2, and 3 GNN layers, respectively, and each one is ubiquitous in GNNs. Figure 4 shows that as the GNN layers increase, EC-LDA performance improves, contrary to Figure 2. This is due to clipping reducing model parameters, which lowers the absolute values and variance of $I_{k}$, resulting in better attack performance. Overall, EC-LDA delivers excellent performance across three kinds of models.
Figure 4: Performance of EC-LDA when the number of GNN layers of the model is 1, 2, and 3, respectively.
Figure 5: EC-LDA’s performance varies with changes in $C$.
Figure 6: The attack performance in EC-LDA, DP-GNN and Label-DP.
Figure 7: The model accuracy in EC-LDA, DP-GNN and Label-DP.
Impact of $C$. In EC-LDA, the server will distributes the clipped model during attack rounds. The method of clipping directly affects the performance of EC-LDA. We employ the $\ell_{2}$ norm clipping method and investigate the impact of $C$ on EC-LDA. Figure 5 shows that the performance of EC-LDA under different $C$. We can observe that as $C$ increases, the performance of EC-LDA deteriorates. This is because with increasing $C$, the clipping intensity decreases. When $C$ exceeds the $\ell_{2}$ norm of the model parameters, the value of $max(1,\frac{\mathbb{N}}{C})$ will be equal to 1, which means no clipping will be applied to the model, leading to a deterioration in EC-LDA’s performance.
Impact of $E$. To study the effect of different $E$ on EC-LDA, we evaluate the performance of EC-LDA under different $E$. We show the experimental results in Table 3. It is evident that EC-LDA performs exceptionally well across varying $E$, with its performance across all datasets almost remaining unaffected by changes in $E$. As introduced in the previous section, we use $\Delta W_{n}^{(FC)}/E$ as an approximation of the gradients for EC-LDA. As $E$ increases, the error between the true gradients and the approximated gradients becomes larger, which results in a slight degradation of the performance of EC-LDA in some cases. Overall, EC-LDA performs effectively across different $E$.
6.4 Defense Performance of Differential Privacy (RQ3)
Differential privacy (DP) is a powerful privacy-preserving technology widely used in machine learning due to its rigorous mathematical definition. DP can defend against various attacks, such as membership inference attacks Hui et al. (2021), adversarial example attacks Lecuyer et al. (2019), and data reconstruction attacks Balle et al. (2022). Specifically, we consider node-level differential privacy DP-GNN Daigavane et al. (2021) and label-level differential privacy Label-DP Ghazi et al. (2021). DP-GNN controls the out-degree of nodes and is a variant of DP-SGD Abadi et al. (2016). Label-DP adds noise to the label matrix to protect the labels of the training dataset.
Figure 6 and Figure 7 show that the performance of EC-LDA and the final accuracy of the FGL model with privacy budget $\epsilon$ on node classification tasks under DP-GNN and Label-DP defenses, respectively. Privacy budget $\epsilon$ controls the magnitude of the added noise: a smaller privacy budget $\epsilon$ results in larger noise, and vice versa. In other words, a smaller privacy budget $\epsilon$ leads to a decline in the model’s final inference performance. From Figure 6, it is evident that the defense effectiveness of Label-DP improves with decreasing privacy budget $\epsilon$. This improvement arises because Label-DP introduces increasing noise to the labels as privacy budget $\epsilon$ decreases. Additionally, the inferred label distribution closely matches the distribution after noise addition due to EC-LDA’s powerfulness in recovering label distributions. This alignment results in a significant disparity between the inferred distributions and actual label distributions, consequently diminishing EC-LDA’s attack effectiveness.
However, as privacy budget $\epsilon$ decreases, the defense effect of DP-GNN remains almost unchanged, which is because the fact that DP-GNN introduces Gaussian noise across all gradient layers, whereas in EC-LDA, the gradients of the last fully connected layer used for attacks are derived from summation along the input dimension. In the summation process, the added Gaussian noise will cancel out the positive and negative, reducing the overall noise, thereby leading to the insensitivity of the defense effect of DP-GNN to privacy budget $\epsilon$. Therefore, the defense methods based on DP-SGD fail to be effective against EC-LDA.
Model Local Epochs Cos-sim JS-div
Cora Facebook Cora Facebook
GCN 1 1.000 1.000 0.000 0.000
3 1.000 1.000 0.000 0.000
5 1.000 1.000 0.001 0.000
GAT 1 1.000 1.000 0.000 0.000
3 1.000 1.000 0.000 0.001
5 1.000 0.999 0.001 0.002
GraphSAGE 1 1.000 1.000 0.000 0.000
3 1.000 1.000 0.000 0.000
5 1.000 1.000 0.000 0.000
Table 3: Attack performance of EC-LDA when $E$ is 1, 3, and 5, respectively.
7 Conclusion
This paper demonstrates the effectiveness of EC-LDA for label privacy in FGL scenarios. We discovered that LDA’s performance is influenced by the variance of node embeddings in GNNs and proposed EC-LDA, which improves attack performance by compressing node embeddings. Extensive experiments on six graph datasets show that EC-LDA outperforms SOTA methods in both node classification and link prediction tasks. We also explored the role of differential privacy in defending against EC-LDA. This paper primarily demonstrates the effectiveness of our attack method. In EC-LDA, the server clips the model’s parameters, actively disrupting the normal training process, and this behavior may be detected. Future work will focus on enhancing the stealthiness of attacks, potentially by training a fishing model with performance and parameter distributions similar to the original model for client deployment, to replace the compressed model.
Appendix A Related Work
Inference Attack against GNNs. Chen et al. [2023] investigated graph data leakage in horizontal federated and vertical federated scenarios. Specifically, link inference attack and attribute inference attack are proposed in the vertical federated scenario, graph reconstruction attack and graph feature attack are proposed in the horizontal federated scenario. Qiu et al. [2022] proposed an inference attack against the relationships between nodes of GNNs in vertical federated scenarios. Meng et al. [2023] inferred the labels of victim nodes by infiltrating the raw graph data, i.e., by adding elaborately designed nodes and edges, and their attacks can only be performed on trained models. However, their attacks focus on the node-level, and there are no extant papers on label distribution inference attacks at the graph-level for FGL.
Label Recovery Attack. The label distribution inference attack is a type of label recovery attack. Zhu et al. [2019] was the first to restore the training sample from gradients. They restored the input data and associated label from the gradients using gradient-matching. This method continuously optimizes the dummy input data and associated label by minimizing the mean square error of the gradients of the dummy sample with respect to the true gradients. Zhao et al. [2020] introduced iDLG and were the first to propose that with a non-negative activation function, privacy label can be extracted with 100% success rate from the signs of gradients at the output layer. Both of these methods are applicable only to single-sample training scenarios. Afterwards, Yin et al. [2021]; Geng et al. [2021] extended the attacks to the mini-batch scenario. Among them, Yin et al. [2021] also requires the use of non-negative activation functions and does not allow duplicate labels within a mini-batch. Therefore, their method is applicable when the number of classes in the dataset is far greater than the batch size. Gu and Bai [2023] believes that the label distribution of the training dataset will leave a footprint in the parameter changes of the model output layer. They use the auxiliary dataset to obtain the connection between the gradients of the output layer and the label distribution of the training data, then train a neural network model accordingly to obtain the label distribution information of the training dataset. Aggarwal et al. [2021] relies on number theory and combinatorics to recover label information from log-loss scores. However, their method is sensitive to the number of classes in the dataset, and as the variety of labels increases, not only does the accuracy decrease, but the inference time also increases. Ma et al. [2023] recovered class-wise embeddings from the gradients and further restored the number of each label. Dang et al. [2021] proposes RLG, which extracts label information with the gradients of the output layer, but RLG requires the use of the soft-max activation function. Additionally, RLG only reveals which labels are used for training, without disclosing the quantity of samples per label. Zhou et al. [2022] proposed PPA, which can infer a client’s label information but can only output the majority class or minority class. It cannot fully reflect the label information. Wainakh et al. [2021] proposed LLG, which utilizes the magnitude and direction of shared gradients to determine whether a specific label is present. LLG has three versions, suitable for different scenarios. Among them, LLG assumes that the attacker can only access shared gradients, LLG* assumes that the attacker can access both the model’s parameters and gradients, and LLG+ assumes that the attacker has an auxiliary dataset. Wang et al. [2019] proposed three attack methods, which can infer whether a specific label appears in the training process, the quantity of each label for a specific client in a round, and the quantity of each label throughout the entire training process, respectively.
Appendix B Experimental Results of the Link Prediction Tasks
Table 4 shows the performance of EC-LDA on all datasets in the link prediction task, where we can see that EC-LDA achieves the optimal values on both metrics, demonstrating an impressive performance.
@@ Appendix B Experimental Results of the Link Prediction Tasks
Table 4: Performance of EC-LDA when training GCN, GAT, and GraphSAGE models on various datasets for link prediction tasks.
@@
Appendix C Label Distribution
We employ a community detection algorithm to partition the graph dataset for each client, which identifies communities within the network(i.e., groups of nodes with high connection densities). Figure 8 shows the distribution of labels assigned to each client by this algorithm for the Cora dataset. From the figure, it is evident that, with a setup of 10 clients, the algorithm simulates a variety of label distributions. Our prior experimental results indicate that EC-LDA consistently delivers strong performance. Therefore, we conclude that EC-LDA can perform well across a range of distributions.
Appendix D Impact of Number of Clients
@@
To investigate the impact of the number of clients on the performance of EC-LDA, we conducted experiments on the Cora dataset, varying the number of clients from 5 to 50. Figure 9 demonstrates the performance of EC-LDA as the number of clients changes. We can see that regardless of the number of clients, the performance of EC-LDA remains excellent, showcasing its wide applicability.
Figure 9: The performance of EC-LDA when the number of clients is 5, 10, 20, and 50, using the WikiCS and Facebook datasets.
Appendix E Datasets and Hyper-parameter Settings
We conduct experiments on six widely used public datasets: Cora Sen et al. [2008], CiteSeer Sen et al. [2008], LastFM Rozemberczki and Sarkar [2020], Facebook Rozemberczki et al. [2021], CoraFull Bojchevski and Günnemann [2017], and WikiCS Mernyei and Cangea [2020]. Cora, CiteSeer, WikiCS, and CoraFull are citation network datasets. Facebook and LastFM are social network datasets. We show the main features of these datasets in Table 5. To demonstrate the effectiveness of EC-LDA under conditions with a large number of labels, we select the CoraFull dataset, which has up to 70 labels.
All datasets use the SGD optimizer. If not specified, all the experiments in this paper use the following setup: all at $E$ of 5, clipping threshold $C$ of 0.01, number of clients of 10, and attack on all the clients in the middle round. All the experiments in this paper are taken with different random seeds to repeat 5 times, and the results are averaged.
Network
Types
Datasets #Nodes #Edges #Features #Classes
Citation Network CiteSeer 2120 7358 3703 6
Cora 2485 10138 1433 7
WikiCS 11311 297033 300 10
CoraFull 18800 125370 8710 70
Social Network Facebook 22470 342004 128 4
LastFM 7624 55612 128 18
Table 5: Main features of the datasets

### v3

@@ IV-3 Extend to Link Prediction
Implementing EC-LDA requires GNN layers with a fully connected final layer and the use of the cross-entropy loss function, which makes EC-LDA particularly versatile within classification tasks on GNNs, such as node classification. Since the link prediction is a binary classification task, EC-LDA is still applicable to link prediction tasks as long as the last fully connected layer of the model has two output units and the cross-entropy loss function is used. For node classification, EC-LDA can attack the label proportions, while for link prediction, it can target the graph density.
V Experiments
@@
  • RQ1 - How effective is EC-LDA under node classification and link prediction tasks with real-world graph datasets?
  • RQ2 - How robust is EC-LDA across different experimental parameters?
  • RQ3 - What is the performance of EC-LDA under differential privacy protection?
V-A Experimental Settings
Datasets: We conduct experiments on six widely used public datasets: Cora [22], CiteSeer [22], LastFM [23], Facebook [24], CoraFull [25], and WikiCS [26]. Cora, CiteSeer, WikiCS, and CoraFull are citation network datasets. Facebook and LastFM are social network datasets. We show the main features of these datasets in Table II. To demonstrate the effectiveness of EC-LDA under conditions with a large number of labels, we select the CoraFull dataset, which has up to 70 labels.
Hyper-parameter Settings: All datasets use the SGD optimizer. If not specified, all the experiments in this paper use the following setup: all at $E$ of 5, clipping threshold $C$ of 0.01, number of clients of 10, and attack on all the clients in the middle round. All the experiments in this paper are taken with different random seeds to repeat 5 times, and the results are averaged.
TABLE II: Statistics of datasets
Types Datasets #Nodes #Edges #Features #Classes
Citation
Network
CiteSeer 2120 7358 3703 6
Cora 2485 10138 1433 7
WikiCS 11311 297033 300 10
CoraFull 18800 125370 8710 70
Social
Network
Facebook 22470 342004 128 4
LastFM 7624 55612 128 18
TABLE III: Performance of EC-LDA and baselines.
@@ V-A Experimental Settings
Evaluation Metrics: Inspired by [27], we use the following two evaluation metrics to fully demonstrate the effectiveness of EC-LDA: cosine similarity (cos-sim) and Jensen-Shannon divergence (JS-div). Cos-sim measures the similarity of two distributions and is applicable when the similarity of vectors is not directly related to the length of the vectors, while JS-div measures the distance between two distributions. The two metrics are calculated using the following two formulas:
$\text{JS-div}(D,D^{*})=\frac{1}{2}\sum_{n=1}^{N}d_{n}\log(\frac{2d_{n}}{d_{n}+d_{n}^{*}})$ (7)
$+\frac{1}{2}\sum_{n=1}^{N}d_{n}^{*}\log(\frac{2d_{n}^{*}}{d_{n}+d_{n}^{*}}),$
$\text{cos-sim}(D,D^{*})=\frac{\sum_{n=1}^{N}(d_{n}d_{n}^{*})}{\sqrt{\sum_{n=1}^{N}d_{n}^{2}}\sqrt{\sum_{n=1}^{N}d_{n}^{*2}}},$ (8)
where $D$ is the inferred label distribution and $D^{*}$ is the ground-truth label distribution. Cos-sim takes values in the range [1,1][-1,1]. The larger the cos-sim, the closer the inferred label distribution is to the ground-truth label distribution, i.e., the more effective the attack is. The value of JS-div ranges from $[0,1]$, and the more effective the attack is, the smaller the JS-div is.
Baselines: We compare EC-LDA with three different attacks which are Infiltrator [28], iLRG [14], and LLG* [12]. Infiltrator infers the label of the victim node by adding a neighbor to the victim node and observing the output of the neighbor. It is worth noting that Infiltrator focuses on node-level attacks, and for comparison, we attack all training nodes with Infiltrator. LLG* and iLRG both reveal the number of each label, and we use the number of samples per label extracted by iLRG and LLG* to compute the label distribution of each client. Similar to EC-LDA, iLRG and LLG* also use the gradients of the last fully connected layer of the model, and when $E$ is greater than 1, we use $\Delta W^{(FC)}/E$ as an approximation of the gradients.
Dummy Training Data: Based on the previous analysis, we know that the smaller the variance of $I$, the better the attack performance. Since the dummy training data is also related to the variance of $I$, we generate it from a Gaussian distribution with a mean of 0 and a standard deviation of 0.001, containing 1000 nodes.
V-B Attack Performance (RQ1)
@@
Node Classification. We evaluate the performance of EC-LDA with all datasets and all GNN types, which, for the node classification task, have a varying number of labels and are widely distributed, with 4 kinds of labels for nodes in the Facebook dataset and 70 kinds of labels for nodes in the CoraFull dataset, the datasets with the smallest and the largest number of classes, respectively. Table III shows the experimental results. From the comparison presented in Table III, it is evident that EC-LDA consistently demonstrates exceptional performance across various datasets and three distinct GNN models, as indicated by the cos-sim scores consistently at or above 0.999 and the JS-div consistently at or below 0.002, aligning closely with the optimal values. Remarkably, EC-LDA’s outstanding performance remains consistent regardless of the number of labels, the number of nodes, and the specific GNN types, showcasing its broad applicability. Additionally, EC-LDA consistently outperforms other methods across all experiments.
TABLE IV: Performance of EC-LDA on link prediction tasks.
@@
Link Prediction. In the link prediction experiments, we set the same number of positive and negative edges. Table IV shows the performance of EC-LDA on all datasets in the link prediction task. In all experiments of link prediction, cos-sim and JS-div reached 1.000 and 0.000, respectively, demonstrating the stunning performance of EC-LDA.
V-C Ablation Experiments (RQ2)
To explore the effectiveness of EC-LDA under varying parameters, we performed extensive ablation experiments on the node classification tasks. Specifically, we explored the impact of the GNN layer depth, local epochs $E$, clipping threshold $C$, number of clients, dummy training data separately. Furthermore, we also demonstrate the label distribution of clients during training.
Impact of the GNN Layer Depth. We illustrate in Fig. 4 the effect of GNN layer depth on the performance of EC-LDA. Specifically, we evaluate three GNN models with 1, 2, and 3 GNN layers, respectively. Fig. 4 shows that as the GNN layers increase, EC-LDA performance improves, contrary to Fig. 2. This is due to clipping reducing model parameters, which lowers the absolute values and variance of $I_{k}$, resulting in better attack performance. Overall, EC-LDA delivers excellent performance across three kinds of models.
Figure 4: Performance of EC-LDA under various the number of GNN layers.
Figure 5: Performance of EC-LDA under various clipping threshold $C$.
@@ V-C Ablation Experiments (RQ2)
Impact of clipping threshold $C$. In EC-LDA, the server will distributes the clipped model during attack rounds. The method of clipping directly affects the performance of EC-LDA. We employ the $\ell_{2}$ norm clipping method and investigate the impact of $C$ on EC-LDA. Fig. 5 shows that the performance of EC-LDA under different $C$. We can observe that as $C$ increases, the performance of EC-LDA deteriorates. This is because with increasing $C$, the clipping intensity decreases. When $C$ exceeds the $\ell_{2}$ norm of the model parameters, the value of $max(1,\frac{\mathbb{N}}{C})$ will be equal to 1, which means no clipping will be applied to the model, leading to a deterioration in EC-LDA’s performance.
Impact of local epochs $E$. In previous experiments we fixed $E$ to be 5, a widely accepted practice in FGL. To study the effect of different $E$ on EC-LDA, we evaluate the performance of EC-LDA under different $E$. We show the experimental results in Table V. It is evident that EC-LDA performs exceptionally well across varying $E$, with its performance across all datasets almost remaining unaffected by changes in $E$. As introduced in the previous section, we use $\Delta W_{n}^{(FC)}/E$ as an approximation of the gradients for EC-LDA. As $E$ increases, the error between the true gradients and the approximated gradients becomes larger, which results in a slight degradation of the performance of EC-LDA in some cases. Overall, EC-LDA performs effectively across different $E$.
Figure 7: Performance of EC-LDA under various client number.
@@
Impact of the Client Number. To investigate the impact of the number of clients on the performance of EC-LDA, we conducted experiments on the Cora dataset, varying the number of clients from 5 to 50. Fig. 7 demonstrates the performance of EC-LDA as the number of clients changes. We can see that regardless of the number of clients, the performance of EC-LDA remains excellent, showcasing its wide applicability.
TABLE V: Performance of EC-LDA across various local epochs $E$.
Model
Local
Epochs
Cos-sim JS-div
Cora Facebook Cora Facebook
GCN 1 1.000 1.000 0.000 0.000
3 1.000 1.000 0.000 0.000
5 1.000 1.000 0.001 0.000
GAT 1 1.000 1.000 0.000 0.000
3 1.000 1.000 0.000 0.001
5 1.000 0.999 0.001 0.002
GraphSAGE 1 1.000 1.000 0.000 0.000
3 1.000 1.000 0.000 0.000
5 1.000 1.000 0.000 0.000
Impact of Dummy Training Data. In EC-LDA, dummy training data is employed to generate parameter $K$, $p_{k,l}$, $I_{k}$, and $\bar{I}$. Table VI illustrates the impact of the standard deviation in dummy training data on the performance of EC-LDA. The results indicate that changes in the distribution of dummy training data have minimal influence on the effectiveness of EC-LDA.
TABLE VI: Performance of EC-LDA under various standard deviation of dummy training data.
Dataset Standard Deviation Cos-sim JS-div
CiteSeer 0.001 1.000 0.001
0.1 1.000 0.001
10 1.000 0.001
Facebook 0.001 1.000 0.000
0.1 1.000 0.000
10 1.000 0.000
Impact of Label Distribution. We employ a community detection algorithm to partition the graph dataset for each client, which identifies communities within the network(i.e., groups of nodes with high connection densities). Fig. 6 shows the distribution of labels assigned to each client by this algorithm for the Cora dataset. From the figure, it is evident that, with a setup of 10 clients, the algorithm simulates a variety of label distributions. Our prior experimental results indicate that EC-LDA consistently delivers strong performance. Therefore, we conclude that EC-LDA can perform well across a range of distributions.
Figure 8: The attack performance of EC-LDA under no defense, DP-GNN defense, and Label-DP defense.
Figure 9: The model accuracy when executing EC-LDA attacks under no defense, DP-GNN defense, and Label-DP defense.
V-D Defense Performance of Differential Privacy (RQ3)
Differential privacy (DP) is a powerful privacy-preserving technology widely used in machine learning due to its rigorous mathematical definition. DP can defend against various attacks, such as membership inference attacks [29], adversarial example attacks [30], and data reconstruction attacks [31]. Specifically, we consider node-level differential privacy DP-GNN [32] and label-level differential privacy Label-DP [33]. DP-GNN controls the out-degree of nodes and is a variant of DP-SGD [34, 35]. Label-DP adds noise to the label matrix to protect the labels of the training dataset. In DP, privacy budget $\epsilon$ controls the scale of the added noise: a smaller privacy budget $\epsilon$ results in larger noise, and vice versa. Fig. 8 and Fig. 9 respectively present the performance of EC-LDA and the node classification accuracy of the FGL model under DP-GNN and Label-DP defenses.
Label-DP. From Fig. 8, it is evident that the defense effectiveness of Label-DP improves with decreasing privacy budget $\epsilon$. Under the Label-DP defense, noise is injected into the label matrix of the training dataset. As the $\epsilon$ value decreases, the magnitude of noise added to the label matrix progressively increases. This results in a widening divergence between the perturbed label matrix and the ground-truth label matrix. Additionally, the inferred label distribution closely matches the distribution after noise addition due to EC-LDA’s powerfulness in recovering label distributions. This alignment results in a significant disparity between the inferred distributions and actual label distributions, consequently diminishing EC-LDA’s attack effectiveness.
DP-GNN. However, as shown in Fig. 8, unlike Label-DP, the defensive efficacy of DP-GNN remains nearly unchanged as $\epsilon$ decreases from 5 to 1. Contrary to the well-established observation that diminishing $\epsilon$ enhances privacy protection, DP-GNN exhibits anomalous $\epsilon$-invariance. This unexpected phenomenon stems directly from the gradient processing mechanics of DP-GNN. As a variant of DP-SGD, DP-GNN first clips the gradients and then injects zero-mean Gaussian noise $\mathcal{N}(0,\sigma^{2})$ into gradients. However, when EC-LDA calculates the label distribution, the gradient it uses is $\Delta W_{l}^{(FC)}$, which is the sum of the gradients of the $l$-th output unit in the last fully connected layer along the input dimension, and $\Delta W_{l}^{(FC)}=\sum_{m}\Delta W_{m,l}^{(FC)}$. We define $\Delta\overline{W}_{l}^{(FC)}$ as the clipped version of gradient $\Delta W_{l}^{(FC)}$, and $\Delta W_{l}^{{}^{\prime}(FC)}$ as the gradient after adding noise to $\Delta\overline{W}_{l}^{(FC)}$. We have:
$\Delta W_{l}^{{}^{\prime}(FC)}$ $=\sum_{m}(\Delta\overline{W}_{m,l}^{(FC)}+\eta_{m,l}),$ (9)
where $\eta_{m,l}\sim\mathcal{N}(0,\sigma^{2})$, then $\mathbb{E}[\eta_{m,l}]=0$, in that way:
$\mathbb{E}[\Delta W_{l}^{{}^{\prime}(FC)}-\Delta\overline{W}_{l}^{(FC)}]$ (10)
$=\mathbb{E}[\sum_{m}\Delta\overline{W}_{m,l}^{(FC)}]+\mathbb{E}[\sum_{m}\eta_{m,l}]-\mathbb{E}[\sum_{m}\Delta\overline{W}_{m,l}^{(FC)}]=0.$
It can be interpreted as: during the summation of $\Delta W_{l}^{{}^{\prime}(FC)}$ over the input dimensions, the Gaussian noise added by DP-GNN undergoes positive-negative cancellation, causing the summed gradient after noise addition, i.e., $\Delta W_{l}^{{}^{\prime}(FC)}$, to approximate the clipped gradient, i.e., $\Delta\overline{W}_{l}^{(FC)}$. Consequently, this leads to the defensive efficacy of DP-GNN being insensitive to $\epsilon$.
VI Related Work
Inference Attack against GNNs. [36] investigated graph data leakage in horizontal federated and vertical federated scenarios. Specifically, link inference attack and attribute inference attack are proposed in the vertical federated scenario, graph reconstruction attack and graph feature attack are proposed in the horizontal federated scenario. [37] proposed an inference attack against the relationships between nodes of GNNs in vertical federated scenarios. [28] inferred the labels of victim nodes by infiltrating the raw graph data, i.e., by adding elaborately designed nodes and edges, and their attacks can only be performed on trained models. However, their attacks focus on the node-level, and there are no extant papers on label distribution inference attacks at the graph-level for FGL.
Label Recovery Attack. The label distribution inference attack is a type of label recovery attack. [7] was the first to restore the training sample from gradients. They restored the input data and associated label from the gradients using gradient-matching. This method continuously optimizes the dummy input data and associated label by minimizing the mean square error of the gradients of the dummy sample with respect to the true gradients. [38] introduced iDLG and were the first to propose that with a non-negative activation function, privacy label can be extracted with 100% success rate from the signs of gradients at the output layer. Both of these methods are applicable only to single-sample training scenarios. Afterwards, [13, 21] extended the attacks to the mini-batch scenario. Extensive research such as [14, 39, 40, 12, 11] demonstrates that gradients can be exploited to extract private label distributions. [41] relies on number theory and combinatorics to recover label information from log-loss scores. However, their method is sensitive to the number of classes in the dataset, and as the variety of labels increases, not only does the accuracy decrease, but the inference time also increases. [42] proposed three attack methods, which can infer whether a specific label appears in the training process, the quantity of each label for a specific client in a round, and the quantity of each label throughout the entire training process, respectively.
VII Limitations
In EC-LDA, the malicious server enhances attack effectiveness by clipping to compress the variance of node embeddings, then distributing this clipped model parameters to clients. This clipping operation leads to an anomalous distributional compression of model parameters. Consequently, a potentially effective defense strategy is for clients to detect potential attacks by identifying these abnormally compressed global model parameters and implementing corresponding countermeasures. However, since EC-LDA only requires a single clipping operation and can be executed at any point during the FGL process, this presents significant challenges for client-side detection and defense. For future work, we will explore effective detection and defense methods against EC-LDA. Additionally, we will concentrate on enhancing the stealthiness of our attacks. To enhance stealth, the server can deploy a “fishing model” that mimics the original model’s performance and parameter distribution while achieving the same effect as embedding compression, allowing the attack to proceed covertly while maintaining its effectiveness.
VIII Conclusion
This paper presents EC-LDA, a label distribution inference attack designed for Federated Graph Learning (FGL). The proposed method was developed by investigating the impact of node embedding variance in Graph Neural Networks (GNNs) on existing LDA approaches. By compressing node embeddings, EC-LDA significantly enhances the attack capability of a malicious server. Extensive experiments on six graph datasets show that EC-LDA outperforms state-of-the-art methods in both node classification and link prediction tasks. Additionally, we demonstrate the robustness of EC-LDA under various parameter settings and under DP protection.
IX ACKNOWLEDGMENTS
This work is supported by Fundamental Research Funds for the Central Universities (Grant No. 40500-20104-222609), National Natural Science Foundation of China Key Program (Grant No. 62132005), Natural Science Foundation of Shanghai (Grant No.22ZR1419100), and CAAI-Huawei MindSpore Open Fund (Grant No.CAAIXSJLJJ-2022-005A).